15.    A certain type of automobile battery is known to have a lifetime which is normally distributed with mean 1,200 days and standard deviation 90 days. For how many days should these batteries be guaranteed if the manufacturer wants to replace only 20% of the batteries sold because they "died" before the guarantee expired?

          A.       1052           B.       1105           C.      1124           D.      1094

16.    A particular television commercial is understood by 30% of first graders and 65% of fourth graders. If one first grader and one fourth grader are randomly selected, what is the probability neither child understands the commercial, assuming their reactions are independent?

  1. 1.05                 B.       0.245          C.      0.95            D.      0.195

17.    The standard deviation of a binomial distribution for which n = 84 and p = 0.22 is:

A.       2.93            B.       8.6              C.      3.8              D.      14.41

18.    A box contains two red and two blue balls. When two balls are randomly selected from that box, without replacement, what is the probability both balls are red?

          A.       1/6              B.       1/4              C.      1/3              D.      1/2
Answer #1

15)

for 20th percentile critical value of z= -0.842
therefore corresponding value=mean+z*std deviation= 1124. days

16)

neither child understands the commercial,=(1-0.3)*(1-0.65)=0.245

17)

standard deviation σ=√(np(1-p))= 3.8

18)

probability both balls are red=(2/4)*(1/3)=1/6